Reduce the quadratic form 5x^2-4xy+8y^2 to its orthogonal canonical form, clearly giving the transformations being used. Also give a rough sketch of the curve representing this canonical form.



1
Expert's answer
2021-02-24T07:07:02-0500

"5x^2 -4xy +8y^2 = 1" (instead of one, any number can be here, you did not specify it)

"(nx \pm my)^2 = c" - canonical form

first of all we should find the square roots of the first and last terms:

"\sqrt{5x^2} = \sqrt{5}*x"

"\sqrt{8y^2} = 2\sqrt{2}*y"


-4 - should be multiplication of "\sqrt{5}" and "2\sqrt{2}" and + free term:


"\sqrt{5}*2\sqrt{2} = 2\sqrt{10}"


Answer: "(\sqrt{5}x + 2\sqrt{2}y)^2 - (4+2\sqrt{10})xy = 1"